For 2A→R , with pure gas A as feed to a PFR ,what is εA=?


A)-0.5


B)0.5


C)2


D)-2

1
Expert's answer
2022-02-17T08:29:06-0500

Gaseous Feed 2A→R

CA0 = 200, CR0 = 0,

CA = 0, CR = 100

VI = 200A + 0R = 200,

VF = 0A + 100R = 100

So εA = (VF - VI) / VI =(100 - 200) / 200 = -0.5

A)-0.5